Timmy Incorporated leases a piece of equipment to Apple Corporation on January 1, 2017.

1. Lease term in years. 4

2. Fair Value of equipment 25,100

3. Book Value of equipment 20,100

4. Lease agreement requires equal annual lease payments, beginning on January 1, 2017 $4,952

Assume accounting periods ends December 31.

5. Estimated economic life of the equipment in years 6

Unguaranteed Residual Value at end of lease term $8100

Expected Residual Value at end of lease term. $8100

6. Lessor and Lessee use straight-line depreciation for all assets.

7. Apple incremental rate of interest. 8%

Timmy implicit rate of interest ( Known to Apple) 5%

8. There is no bargain purchase option, ownership of the lease does not transfer at the end of the lease term, and the asset is not of a specialized nature.

(a) Does the lease pass the Present Value test?

(b) Determine the Lease Liability at commencement of the lease.

(c) Calculate depreciation expense for the lessor.
Answer #1

a) Present value test for lease is a criteria used by the lessee to account for the leased property. Under this criteria, if the present value of minimum lease payments is equal to or greater than 90% of the fair market value of the leased asset then it is considered as capital lease. To check whether the given lease pass the present value test or not, we need to calculate the present value of minimum lease payments which is shown as follows:-

The discounting rate to be used = Apple incremental rate of interest = 8%

Lease Period = 4 years

Present Value of Minimum Lease Payments = PV of annual lease payments+PV of guaranteed Residual value

PV of Annual Lease Payments = Annual Lease Payments*PVAF(8%, 3 yrs)+First lease Payment

= ($4,952*2.57710)+$4,952 = $17,714

(As the lease payments are paid at the beginning of each year, first lease payment is not discounted)

(PV of guaranteed residual value is zero as guaranteed residual value is zero.)

Present Value of Minimum Lease Payments = $17,714+$0 = $17,714

Fair Market Value of Leased Asset = $25,100

90% of Fair Market Value = $25,100*90% = $22,590.

As present value of minimum lease payments ($17,714) is less than 90% of fair market value of leased asset ($25,100), the given lease does not pass the present value test.

b) Lease liability at the commencement of lease will equal to present value of minimum lease payments i.e. $17,714 (as calculated in part 1).

c) Depreciation expense for the lessor as per straight line method is calculated as follows:-

Depreciation Expense per year = (Book Value of Equipment/Estimated Economic life)

= $20,100/6 years = $3,350 per year